# Fourier Analysis of Finite Difference Schemes for the Helmholtz Equation: Sharp Estimates and Relative Errors

### v1

@@ 3.5 Relative error estimates of the classical 3-point centred scheme
Suppose $k>\frac{4\pi}{\pi-\sqrt{16-\pi^{2}}}\approx 18.88$, $10<\frac{kh}{2}\leq C_{\mu}<$ 1, and there exists a sequence of $h=\frac{1}{N}$ going to zero such that
$\tilde{\sigma}_{k}:=\min_{\xi,h}\left|k-\frac{2}{h}\sin\frac{\xi h}{2}\right|>0\;\text{over $\xi\in\{1,\ldots,N-1\}\pi$ and the sequence of $h$ going to zero}.$
@@
Recall from Remark 1 that as $N\to\infty$ we have $f_{\mathrm{high}}\to 0$ and $f_{\mathrm{low}}\to f$ for a fixed $f\in H^{p}_{0}(0,1)$. So the relative errors converge faster than the displayed powers of $h$.
@@
Suppose $k>\frac{4\pi}{\pi-\sqrt{16-\pi^{2}}}\approx 18.88$, $10<\frac{kh}{2}\leq C_{\mu}<$ 1, and there exists a sequence of $h=\frac{1}{N}$ going to zero such that
$\tilde{\sigma}_{k}:=\min_{\xi,h}\left|k-\frac{2}{h}\sin\frac{\xi h}{2}\right|>0\;\text{over $\xi\in\{1,\ldots,N-1\}\pi$ and the sequence of $h$ going to zero}.$
@@
which holds with equality if and only if $\hat{f}(\xi)=0$ for $\xi\in\{\pi,\ldots,(N-1)\pi\}\backslash\{\xi_{*}\}$ where $\{\xi_{*}\}$ is the maximizer set of $|\lambda-\lambda^{h}|/|\xi^{p}\lambda^{h}|$ over $\xi\in\{1,\ldots,N-1\}\pi$. So, if $\xi_{*}$ does not change with $N$, then the equality holds for the fixed $f(x)=\sin(\xi_{*}x)$ and $|u_{\mathrm{low}}|_{p}/\|u\|=\xi_{*}^{p}$. Our intention is to find, for a given wavenumber $k$, a source $f$ independent of the mesh size $h=1/N$ (but $f$ may depend on $k$) such that the upper bound is attainable. We will see that $p=2$ is an appropriate choice.
@@
$\tilde{\sigma}_{k}:=\min_{\xi,h}\left|k-\frac{2}{h}\sin\frac{\xi h}{2}\right|>0\;\text{over $\xi\in\{1,\ldots,N-1\}\pi$ and the sequence of $h$ going to zero}.$
@@
$\phi_{rel}(\theta)=\frac{\theta^{2}-\sin^{2}\theta}{\theta^{2}(\sin^{2}\theta-\mu^{2})}\quad\text{is convex in }\theta\text{ for }0<\theta_{k}:=\frac{h}{2}k^{h}=\arcsin\mu<\theta<\frac{\pi}{2}.$
Indeed, denoting by $s:=\sin\theta$, $c:=\cos\theta$, after some calculations we find
$\phi^{\prime\prime}_{rel}(\theta)=\frac{\theta^{2}\left(\theta^{2}-\mu^{2}\right)(1-s^{2})(2\mu^{2}+6s^{2})+2s^{2}\left(s^{2}-\mu^{2}\right)\left(-\mu^{2}(-3+\theta^{2})+\theta^{4}-3s^{2}\right)-8\mu^{2}\theta sc(s^{2}-\mu^{2})}{\theta^{4}\left(s^{2}-\mu^{2}\right)^{3}}.$
It suffices to show that the numerator of $\phi_{rel}^{\prime\prime}(\theta)$ is positive. Note that for $0<\theta<\frac{\pi}{2}$, we have $c<1-\frac{1}{2}\theta s$, $\theta-\frac{1}{6}\theta^{3}<s<\theta-\frac{1}{6}\theta^{3}+\frac{1}{120}\theta^{5}$, and $10<\mu<s<$ 1. Combining these inequalities, we used the Mathematica command Resolve to find the numerator of $\phi^{\prime\prime}_{rel}(\theta)$ is indeed positive. So $\phi_{rel}$ is convex when $\sin\theta>\mu$.
@@
$\tilde{\sigma}_{k}:=\min_{\xi,h}\left|k-\frac{2}{h}\sin\frac{\xi h}{2}\right|>0\;\text{over $\xi\in\{1,\ldots,N-1\}\pi$ and the sequence of $h$ going to zero}.$
@@ 4 Visual analysis of dispersion correction schemes
Then, the discretization error of the new scheme vanishes when $f=0$. An error analysis when $f\neq 0$ is given by [8] using the framework ‘consistency + stability $\Rightarrow$ convergence’, which shows the maximum error is bounded from above by $(\frac{1}{k}\|f^{\prime\prime}\|_{L^{\infty}}+k^{2}\|f\|_{L^{\infty}})h^{2}$ up to a constant factor.
@@
To evaluate the two schemes (36) and (37) more precisely, we use the Fourier analysis as a visual tool. Since the downsampling and aliasing errors are smaller (for smooth $f$) than the operator error, we will focus on the latter. The operator symbols of the left hand sides of (36) and (37) are
$\lambda^{h}_{k}:=\tilde{k}^{2}-\frac{4}{h^{2}}\sin^{2}\frac{\xi h}{2},\quad\lambda_{\Delta}^{h}:=k^{2}-\frac{k^{2}}{\tilde{k}^{2}}\frac{4}{h^{2}}\sin^{2}\frac{\xi h}{2},\quad\xi\in\{1,\ldots,N-1\}\pi\text{ for }h=\frac{1}{N}.$ (38)
@@
where $\lambda^{h}$ is the discrete symbol. For the classical centred scheme, $\lambda^{h}=k^{2}-\frac{4}{h^{2}}\sin^{2}\frac{\xi h}{2}$. For the scheme (36) or (37), it is $\lambda^{h}_{k}$ or $\lambda_{\Delta}^{h}$ in (38). From (28), (30), (34) and (35), it is essentially the symbol error $\psi$, $\psi_{e}$ or $\psi_{rel}$ that determines the convergence in terms of absolute/relative error in the $L^{2}$- or $H^{1}$-semi-norm.
Figure 3 shows the symbol errors $\psi_{e}$ (39) for the three schemes under $h$-refinement. We see that the dispersion correction schemes have lower maxima of $\psi_{e}$ than the classical one, at the price of larger $\psi_{e}$ at smaller $\xi$ (roughly $<\frac{5}{6}k$ or $<\frac{2}{3}k$). We also see that the scheme (37) modifying the discrete Laplacian is better than the scheme (36) modifying the wavenumber. For all the three schemes, $\max\psi_{e}$ is always attained around the wavenumber $k$, and the order of $\max\psi_{e}$ is $h^{2}$, which corroborates Lemma 11.
Figure 3: Symbol errors $\psi_{e}$ (39) for the classical (8) & dispersion correction schemes (36), (37): $h$-refinement
Next, the influence of the wavenumber $k$ is demonstrated, see Figure 4 in which $k$ is quadrupled while $kh$ is halved. This clearly shows that the classical scheme can not converge in this setting with $k^{3}h^{2}$ fixed. This corresponds to the order $k^{3}h^{2}$ proved in Lemma 10. Figure 4 shows also that the dispersion correction schemes have the symbol error of order $k^{2}h^{2}$ which translates to the same order of $|u-u^{h}|_{1}$ by (28).
Figure 4: Symbol errors $\psi$ (39) for the classical (8) & dispersion correction schemes (36), (37): $kh$-refinement
Finally, the relative symbol error $\psi_{rel}$ is visualized in Figure 5 where $kh$ is fixed while $k$ is doubled. According to (35), $\max\psi_{rel}$ has to be multiplied with $|u_{\mathrm{low}}|_{3}/|u|_{1}$ to give the upper bound of $|u-u^{h}|_{1}/|u|_{1}$, and when the upper bound is attained, $|u_{\mathrm{low}}|_{3}/|u|_{1}$ may contribute some powers of $k$ (see Theorem 4). For the classical scheme, we see from Figure 5 that $\psi_{rel}$ is of order $kh^{2}$ which corroborates Lemma 14. For the dispersion correction scheme (36) modifying $k$, the relative symbol error $\psi_{rel}$ does not decrease with $h$ (while $kh$ is fixed) at the first frequency $\xi=\pi$, which slows down also the convergence at the other low frequencies. The dispersion correction scheme (37) modifying the discrete Laplacian has an almost equidistributed $\psi_{rel}$ over $\xi$ as seen in the figure, and $\max\psi_{rel}$ is attained at $\xi=(N-1)\pi$ (order $k$) and is of order $h^{2}$. So the relative error $|u-u^{h}|_{1}/|u|_{1}$ of the scheme (37) is of order $k^{2}h^{2}$.
Figure 5: Symbol errors $\psi_{rel}$ (39) for the classical (8) & dispersion correction schemes (36), (37): $kh$ fixed
@@ 5 Discussions
There are some limitations of the proposed approach in this work. First, it only works on uniform grids of simple geometry. Second, with radiation boundary conditions the generalized Fourier decomposition based on the Sturm-Liouville problem will lead to a non-orthogonal basis, and the current approach does not work directly.
For future work, we will consider models in 2D/3D and analyze some compact high-order schemes. Moreover, it may be worthwhile to try optimizing discretization parameters based on the symbol errors.

### v2

@@ 3.5 Relative error estimates of the classical 3-point centred scheme
Suppose $k>\frac{4\pi}{\pi-\sqrt{16-\pi^{2}}}\approx 18.88$, $0<\frac{kh}{2}\leq C_{\mu}<1$, and there exists a sequence of $h=\frac{1}{N}$ going to zero such that
[…]
@@
Recall from Remark 1 that as $N\to\infty$ we have $f_{\mathrm{high}}\to 0$ and $f_{\mathrm{low}}\to f$ for a fixed $f\in H^{p}_{0}(0,1)$. So the relative errors converge faster than the displayed powers of $h$. But if $|f_{\mathrm{high}}|_{p}=O(k^{p})$ and $kh=O(1)$, then Lemma 12 says the relative error is bounded in $L^{2}$ and $H^{1}$.
@@
Suppose $k>\frac{4\pi}{\pi-\sqrt{16-\pi^{2}}}\approx 18.88$, $0<\frac{kh}{2}\leq C_{\mu}<1$, and there exists a sequence of $h=\frac{1}{N}$ going to zero such that
[…]
@@
[…]
@@
[…]
@@
$\phi_{rel}(\theta)=\frac{\theta^{2}-\sin^{2}\theta}{\theta^{2}(\sin^{2}\theta-\mu^{2})}\quad{\color[rgb]{1,0,0}\text{has no local maximum for }0<\theta_{k}:=\frac{h}{2}k^{h}=\arcsin\mu<\theta<\frac{\pi}{2}}.$
Indeed, we calculate
$\theta^{3}\left(\sin^{2}\theta-\mu^{2}\right)^{2}\phi_{rel}^{\prime}(\theta)=-2\left(\theta^{3}\sin\theta\cos\theta-\sin^{4}\theta+\mu^{2}\sin^{2}\theta-\theta\mu^{2}\sin\theta\cos\theta\right)$
and a criticial point $\theta_{0}\in(\arcsin\mu,\frac{\pi}{2})$ of $\phi_{rel}$ is a root of
$\mu^{2}=\frac{\theta^{3}\cos\theta-\sin^{3}\theta}{\theta\cos\theta-\sin\theta}=:v(\theta).$ (39)
By calculations, we find
$(\theta\cos\theta-\sin\theta)^{2}v^{\prime}(\theta)=(\theta-\sin\theta)(\theta+\sin\theta)\left(\theta\sin^{2}\theta+3\theta\cos^{2}\theta-3\sin\theta\cos\theta\right)>0,$
because $\left(\theta\sin^{2}\theta+3\theta\cos^{2}\theta-3\sin\theta\cos\theta\right)^{\prime}=4\sin^{2}\theta-4\theta\sin\theta\cos\theta>0$ for $\theta\in(0,\frac{\pi}{2})$. Hence, $v^{\prime}(\theta)>0$ and there is at most one critical point of $\phi_{rel}$. Suppose $\theta_{0}\in(\arcsin\mu,\frac{\pi}{2})$ is a critical point of $\phi_{rel}$. By substitution with (39), we find
$\theta_{0}^{5}\left(\theta_{0}^{2}-\sin^{2}\theta_{0}\right)\phi_{rel}^{\prime\prime}(\theta_{0})=2\sin\theta_{0}\cos\theta_{0}(\tan\theta_{0}-\theta_{0})\left(3\theta_{0}+\theta_{0}\tan^{2}\theta_{0}-3\tan\theta_{0}\right)>0,$
because $\left(3\theta+\theta\tan^{2}\theta-3\tan\theta\right)^{\prime}=2\tan\theta\left(\theta\sec^{2}\theta-\tan\theta\right)>0$ for $\theta\in(0,\frac{\pi}{2})$. That means $\theta_{0}$ can only be a local minimum.
@@
[…]
@@ 4 Visual analysis of dispersion correction schemes
Then, the discretization error of the new scheme vanishes when $f=0$. An error analysis when $f\neq 0$ is given by [9] using the framework “consistency + stability $\Rightarrow$ convergence”, which shows the maximum error is bounded from above by $(\frac{1}{k}\|f^{\prime\prime}\|_{L^{\infty}}+k^{2}\|f\|_{L^{\infty}})h^{2}$ up to a constant factor.
@@
As we have seen, the 1D dispersion correction is determined up to a constant factor. If we follow a third approach proposed by [40] originally for 2D and 3D problems, then we would get
$\left(k^{2}\!-\!\frac{k^{2}}{\tilde{k}^{2}}\frac{2}{h^{2}}\right)u^{h}(x_{j})\!+\!\frac{k^{2}}{\tilde{k}^{2}}\frac{1}{h^{2}}u^{h}(x_{j-1})\!+\!\frac{k^{2}}{\tilde{k}^{2}}\frac{1}{h^{2}}u^{h}(x_{j+1})\!=\!\frac{kh}{2}\cot\frac{kh}{2}f(x_{j}),\;u^{h}(x_{0})\!=\!g_{0},\;u^{h}(x_{N})\!=\!g_{1}.$ (42)
To evaluate the HZthree dispersion free schemes (40), (41) and (42), we use the Fourier analysis as a visual tool. Since the downsampling and aliasing errors are smaller (for smooth $f$) than the operator error, we will focus on the latter. The operator symbols of (40), (41) and (42), that map $u^{h}$ to $f^{h}$, are
$\lambda^{h}_{k}\!:=\!\tilde{k}^{2}\!-\!\frac{4}{h^{2}}\sin^{2}\frac{\xi h}{2},\;\;\lambda_{\Delta}^{h}\!:=\!k^{2}\!-\!\frac{k^{2}}{\tilde{k}^{2}}\frac{4}{h^{2}}\sin^{2}\frac{\xi h}{2},\;\;{\color[rgb]{1,0,0}\lambda_{\Delta,f}^{h}\!:=\!\frac{2}{kh}\tan\frac{kh}{2}\lambda_{\Delta}^{h},}\;\;\xi\!\in\!\{1,\ldots,N\!-\!1\}\pi\text{ for }h\!=\!\frac{1}{N}.$ (43)
@@
where $\lambda^{h}$ is the discrete symbol. For the classical centred scheme, $\lambda^{h}=k^{2}-\frac{4}{h^{2}}\sin^{2}\frac{\xi h}{2}$. For the scheme (40), (41) or (42), it is $\lambda^{h}_{k}$, $\lambda_{\Delta}^{h}$ or $\lambda_{\Delta,f}^{h}$ in (43). From (29), (32), (37) and (38), it is essentially the symbol error $\psi$, $\psi_{e}$ or $\psi_{rel}$ that determines the convergence in terms of absolute/relative error in the $L^{2}$- or $H^{1}$-semi-norm.
Figure 3 shows the symbol errors $\psi_{e}$ (44) for the four schemes under $h$-refinement. We see that the dispersion free schemes have lower maxima of $\psi_{e}$ than the classical one, at the price of larger $\psi_{e}$ at smaller $\xi$ (roughly $\frac{5}{6}k$ or $\frac{2}{3}k$). We also see that the scheme (42) modifying both the discrete Laplacian and the source performs the best. The order of $\max\psi_{e}$ is $h^{2}$, which corroborates Lemma 11.
Figure 3: Symbol errors $\psi_{e}$ (44) for the classical (9) & dispersion free (40), (41), (42): $h$-refinement
Next, the influence of the wavenumber $k$ is demonstrated, see Figure 4 in which $k$ is quadrupled while $kh$ is halved. This clearly shows that the classical scheme can not converge in this setting with $k^{3}h^{2}$ fixed. This corresponds to the order $k^{3}h^{2}$ proved in Lemma 10. Figure 4 shows also that the schemes (40) and (41) have the symbol error of order $k^{2}h^{2}$ which translates to the same order of $|u-u^{h}|_{1}$ by (29), while the scheme (42) is even better and of  [^2] This is because the order in $h$ for fixed $k$ must be 2, and it is observed that the scaling $k\to 4k$, $h\to h/8$ makes the error scaled by $4^{3}/8^{3}=2^{-3}$, so that $4^{\alpha}/8^{2}=2^{-3}$ gives the right exponent $\alpha=\frac{3}{2}$ for $k$. order $k^{3/2}h^{2}$.
Figure 4: Symbol errors $\psi$ (44) for the classical (9) & dispersion free (40), (41), (42): khkh-refinement
Finally, the relative symbol error $\psi_{rel}$ is visualized in Figure 5 where $kh$ is fixed while $k$ is doubled. According to (38), $\max\psi_{rel}$ has to be multiplied with $|u_{\mathrm{low}}|_{3}/|u|_{1}$ to give the upper bound of $|u-u^{h}|_{1}/|u|_{1}$, and when the upper bound is attained, $|u_{\mathrm{low}}|_{3}/|u|_{1}$ may contribute some powers of $k$ (see Theorem 4). For the classical scheme, we see from Figure 5 that $\psi_{rel}$ is of order $kh^{2}$ which corroborates Lemma 14. For the dispersion free schemes (40) and (42), the relative symbol error $\psi_{rel}$ does not decrease with $h$ (while $kh$ is fixed) at the first frequency $\xi=\pi$, which slows down also the convergence at the other low frequencies. The dispersion free scheme (41) modifying the discrete Laplacian has an almost equidistributed $\psi_{rel}$ over $\xi$ as seen in the figure, and $\max\psi_{rel}$ is attained at $\xi=(N-1)\pi$ (order $k$) and is of order $h^{2}$. So the relative error $|u-u^{h}|_{1}/|u|_{1}$ of the scheme (41) is of order $k^{2}h^{2}$.
Figure 5: Symbol errors $\psi_{rel}$ (44) for the classical (9) & dispersion free (40), (41), (42): khkh fixed
@@ 5 Discussions
There are some limitations of the proposed approach in this work. First, it only works on uniform grids of simple geometry. Second, with radiation boundary conditions the generalized Fourier decomposition based on the Sturm-Liouville problem will lead to a non-orthogonal basis, and the current approach does not work directly. It might be interesting to try the singular value decomposition i.e. considering the least-squared form of the Sturm-Liouville problem, which diagonalises the Helmholtz operator with two orthogonal bases.
However, the proposed approach provides a new tool for quantitatively evaluating linear schemes, in complement to dispersion analysis. The latter considers only the zero source case and does not lead to error estimates.
For future work, we will consider models in 2D/3D with simple geometries and analyze some compact high-order schemes, which is more challenging e.g. in finding the maximum of a multivariable symbol. Moreover, it may be worthwhile to try optimizing discretization parameters based on the symbol errors.
It is still an open problem to give wavenumber explicit error estimates of finite difference schemes actually used in applications for variable coefficient Helmholtz problems. The variational theory of FDMs developed in [28] may be useful.